# Rational motives on pro-algebraic stacks

### v3

@@ Acknowledgement
This project started after a talk with Torsten Wedhorn about motives on limits of algebraic stacks and I would like to thank him for various discussions and his interest in this project. Further, I would like to thank Rizacan Ciloglu, David G. Davis, Martin Gallauer, Benjamin Hennion, Ryomei Iwasa, Hyungseop Kim, Eike Lau, Timo Richarz, Jakob Scholbach and Thibaud van den Hove for helpful discussions.
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - project number 524431573, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124 and by the LOEWE grant ‘Uniformized Structures in Algebra and Geometry’.
@@ 3. $\bullet$
    The $\infty$-category of small $\infty$-groupoids is denoted by $SS\SS$.
@@ 5. $\bullet$
    The $\infty$-category of presentable stable $\infty$-categories with continuous functors, i.e. filtered colimit preserving functors, is denoted by $\textup{Pr}^{\textup{L}}_{\textup{st}}$.
@@ 7. $\bullet$
Throughout, let $B$ be an excellent noetherian scheme of dimension $\leq 2$ and let $S$ be a regular scheme of finite type over $B$.
@@ 1. $\bullet$
    An Artin stack is an algebraic stack in the sense of [Sta22].
@@ 2. $\bullet$
    Let $X$ be an $S$-scheme of finite type. By $\textup{DM}(X)$, we denote the DG-category of Beilinson-motives (cf. [CD19, §14]). Roughly, these are $\mathbb{A}^{1}$-local étale sheaves on $X$ with values in ${\mathcal{D}}(\mathbb{Q})$ such that $\mathbb{P}^{1}$ is $\otimes$-invertible.
2 (Motivic) six functor formalisms on pro-algebraic stacks
@@
In this article, we want to extend the 6-functor formalism to limits of Artin stacks, for example the stack of Barsotti-Tate groups. The stack of such groups can be obtained as a limit of its truncations. We will explain later that these truncation maps are given by affine bundles (see also [Bro18]). In particular, if we have a 6-functor formalism for motives on such stacks, we can analyze their motive by truncating. All of this will be made more precise later on. But first let us establish a general theory on inverse systems of Artin stacks that may have non-representable transition maps.
@@
Before we begin with our main constructions and definitions, we want to recall the theory of rational motives on Artin stacks. We will follow the construction of Richarz-Scholbach to obtain the DG-category of rational motives on prestacks (cf. [RS20, Def. 2.2.1]). Later on we will specialize this to our case. In the following, we will omit some details and refer to loc. cit. for an explicit construction.
@@
We denote by $\textup{AffSch}^{\textup{ft}}_{S}$ the category of affine schemes over $S$. As $\textup{AffSch}^{\textup{ft}}_{S}$ is essentially small we replace $\textup{AffSch}^{\textup{ft}}_{S}$ by a small skeleton of objects of interest in this article. We denote by $\textup{AffSch}^{\kappa}_{S}$ the pro-$\kappa$-completion of $\textup{AffSch}^{\textup{ft}}_{S}$. Again, this category is small (cf. [KS06, Lem. 6.1.2]).
@@
A prestack over $S$ is a functor from $\textup{AffSch}^{\kappa}_{S}$ to $\infty$-groupoids and we denote the $\infty$-category of prestacks over $S$ by $\textup{PreStk}_{S}$.
@@
The functor $\textup{DM}\colon X\mapsto\textup{DM}(X)$ denotes the functor of Beilinson motives on finite type $S$-schemes. We extend DM to prestacks in the following way. First we left Kan extend DM from $\textup{AffSch}^{\textup{ft}}_{S}$ to $\textup{AffSch}^{\kappa}_{S}$ and then right Kan extend along the Yoneda embedding $\textup{AffSch}^{\kappa}_{S}\hookrightarrow\textup{PreStk}_{S}$, where the transition maps are given by *-pullbacks. This construction yields a functor
$\textup{DM}^{*}\colon\textup{PreStk}_{S}\rightarrow\textup{DGCat}_{\textup{cont}}.$
@@
If instead of *-pullback, we use the !!!-pullback, we denote this functor by $\textup{DM}^{!}$. However, the restriction to schemes locally of finite type over $S$ yields equivalent descriptions of $S$. This is due to the fact that in our setting DM for schemes satisfies h-descent (cf. [CD19]). Thus, both $\textup{DM}^{!}$ and $\textup{DM}^{*}$ can be computed via descent. This follows as locally finite type schemes admit an open covering by affines of finite type and !!!- resp. *-pullback agree up to shift and twist. Note that the same argumentation can be used to see that $\textup{DM}^{*}(X)\simeq\textup{DM}^{!}(X)$ for any algebraic space, as the diagonal of an algebraic space is representable by a scheme.
@@
Using étale descent, we can glue the 6-functor formalism of DM to the restriction of $\textup{DM}^{*}$ to algebraic spaces, as in [Kha19] or equivalently by using the DESCENT algorithm of Liu-Zheng (cf. [LZ17]). Similar constructions can also be found in [RS20], where ind-Artin stacks and ind-schemes are considered.
@@
Finally, we can now apply the same arguments to glue the 6-functor formalism to the restriction of $\textup{DM}^{*}$ to Artin stacks locally of finite type. As before the existence of a smooth cover on an Artin stack shows $\textup{DM}^{!}(X)\simeq\textup{DM}^{*}(X)$ for any Artin stack $X$ locally of finite type over $S$.
To ease notation, we will from here on denote by DM the extension $\textup{DM}^{*}$. For the !!!-version, we will keep the notation $\textup{DM}^{!}$.
@@
Let $x\colon{\mathcal{I}}^{\operatorname{op}}\rightarrow\textup{PreStk}_{S}$ be a filtered system of prestacks and let $X\coloneqq\lim x$ be the associated limit prestack. We call the tuple $(x,X)$ a $-{\mathcal{I}}$-algebraic stack, if $x(i)$ is an algebraic stack locally of finite type over $S$ for all $i\in{\mathcal{I}}$.
If $(x,X)$ is a pro-algebraic stack, then we denote by $X_{i}\coloneqq x(i)$ and the transition maps $X_{j}\rightarrow X_{i}$ by $x_{ij}$, for $i\rightarrow j\in{\mathcal{I}}$.
A morphism $f\colon(x,X)\rightarrow(y,Y)$ of pro-algebraic stacks is a morphism $f_{\bullet}\colon x\rightarrow y$. If ${\bf P}$ is a property of a morphism of algebraic stacks, we will say that $f$ has property ${\bf P}$ if each $f_{i}\colon X_{i}\rightarrow Y_{i}$ has property ${\bf P}$.
We denote by $\textup{ProAlg}^{{\mathcal{I}}}$ the $\infty$-category of $-{\mathcal{I}}$-algebraic stacks.
If $X$ is an algebraic stack, we will abuse notation to denote the constant diagram $n\mapsto X$ (here the transition maps are given by $\operatorname{id}_{X}$) with $X$ itself.
@@
If $f\colon(x,X)\rightarrow(y,Y)$ is a morphism of $-{\mathcal{I}}$-algebraic stacks, then we call $f$ cartesian if each square
[…]
is a pullback diagram. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks adjointable if the morphism $y_{ij}^{*}f_{i*}\rightarrow f_{j*}x_{ij}^{*}$ induced by each square above is an equivalence.
Example 2.4.
Let us give two important examples of adjointable morphisms of $-{\mathcal{I}}$-algebraic stacks.
  1. (1)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a cartesian morphism of monoidal $-{\mathcal{I}}$-algebraic stacks, then by smooth base change $f$ is adjointable.
  2. (2)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of $-{\mathcal{I}}$-algebraic stacks. Assume that each $y_{ij*}$ is fully faithful (e.g. when $Y$ is equivalent to a constant diagram of induced by an algebraic stack) and $(x,X)$ is strict, then simple computation shows that $f$ is adjointable.
Definition 2.5.
Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack. Then we say that $(x,X)$ is
@@
Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack and assume that each $x_{ij}$ is an affine bundle for all $i\rightarrow j\in{\mathcal{I}}$, i.e. Zariski locally isomorphic to an affine space, then $(x,X)$ is a classical monoidal strict $-{\mathcal{I}}$-algebraic stack.
@@
[…]
The above can be generalized to the following case. We can replace $\operatorname{Spec}(k)$ by $S$ and assume that $S$ is connected. Let $\bar{s}$ be a geometric point of $S$. Then we can define $F_{S}$ as the category of pointed étale covers over $(S,\bar{s})$. The construction above can be repeated in this case for any smooth Artin stack ${\mathfrak{X}}$ over $S$.
@@
Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack. Then we define the $\infty$-category of rational motives on $(x,X)$ as

### v4

@@ Acknowledgement
This project started after a talk with Torsten Wedhorn about motives on limits of algebraic stacks and I would like to thank him for various discussions and his interest in this project. I want to thank Chirantan Chowdhury for his comments about correspondences and the idea to use the formalism of bisimplicial sets to extend diagrams to correspondences. I also thank the anonymous referee for all their suggestions. Further, I would like to thank Rızacan Çiloğlu, Martin Gallauer, Benjamin Hennion, Ryomei Iwasa, Hyungseop Kim, Timo Richarz, Jakob Scholbach for helpful discussions.
This project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - project number 524431573, the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124.
@@ 3. $\bullet$
    The $\infty$-category of small $\infty$-groupoids is denoted by $S\SS$.
@@ 5. $\bullet$
    The $\infty$-category of presentable $\infty$-categories with continuous functors, i.e. colimit preserving functors, is denoted by $\Pr^{L}$.
  6. $\bullet$
    We denote by $\textup{Pr}^{L,\otimes}$ the $\infty$-operad, whose underlying $\infty$-category is $\Pr^{L}$.
  7. $\bullet$
    The $\infty$-category of presentable stable $\infty$-categories with continuous functors is denoted by $\textup{Pr}^{\textup{L}}_{\textup{st}}\coloneqq\mathrm{Mod}_{\operatorname{Sp}}(\textup{Pr}^{L,\otimes})$.
@@ 9. $\bullet$
Throughout, let $S$ be an excellent noetherian scheme of finite dimension.
@@ 1. $\bullet$
    An algebraic stack is in the sense of [Sta25], also referred to as Artin stack.
@@ 2. $\bullet$
    Let $X$ be an $S$-scheme of finite type. By $\textup{DM}(X)$, we denote the DG-category of Beilinson-motives [CD19, §14]. Roughly, these are $\mathbb{A}^{1}$-local étale sheaves on $X$ with values in ${\mathcal{D}}(\mathbb{Q})$ such that $\mathbb{P}^{1}$ is $\otimes$-invertible.
2 Motives on pro-algebraic stacks
@@
In this article, we aim to extend the six-functor formalism to limits of algebraic stacks, such as the stack of Barsotti–Tate groups. This stack can be realized as a limit of its truncations. We develop a general theory for inverse systems of algebraic stacks, allowing for non-representable transition maps.
@@
Before introducing our main constructions and definitions, we recall the theory of rational motives on algebraic stacks. Following the construction of Richarz–Scholbach, we obtain the DG-category of rational motives on prestacks as defined in [RS20, Def. 2.2.1]. We will later specialize this to our specific context. In what follows, we omit some details and refer the reader to loc. cit. for an explicit construction.
@@
Let $\textup{AffSch}^{\textup{ft}}_{S}$ denote the category of affine schemes of finite type over $S$. Since $\textup{AffSch}^{\textup{ft}}_{S}$ is essentially small, we replace it with a small skeleton consisting of the relevant objects for our purposes. We write $\textup{AffSch}^{\kappa}_{S}$ for the pro-$\kappa$-completion of $\textup{AffSch}^{\textup{ft}}_{S}$, which is again small by [KS06, Lem. 6.1.2].
@@
A prestack over $S$ is a functor from $\textup{AffSch}^{\kappa}_{S}$ to $\infty$-groupoids. The $\infty$-category of prestacks over $S$ is denoted by $\textup{PreStk}_{S}$.
@@
The functor $\textup{DM}\colon X\mapsto\textup{DM}(X)$ denotes the functor of Beilinson motives on finite type $S$-schemes. We extend DM to prestacks as follows. First, we perform a left Kan extension of DM from $\textup{AffSch}^{\textup{ft}}_{S}$ to $\textup{AffSch}^{\kappa}_{S}$, followed by a right Kan extension along the Yoneda embedding $\textup{AffSch}^{\kappa}_{S}\hookrightarrow\textup{PreStk}_{S}$. The transition maps in this process are given by *-pullbacks. This yields a functor
$\textup{DM}^{*}\colon\textup{PreStk}_{S}^{\operatorname{op}}\to\textup{DGCat}_{\textup{cont}}.$
@@
If we instead use !!-pullbacks in place of *-pullbacks, we denote the resulting functor by $\textup{DM}^{!}$. However, when restricted to schemes locally of finite type over $S$, these two constructions are equivalent. This equivalence follows from the fact that, in our setting, DM satisfies h-descent on schemes [CD19]. Consequently, both $\textup{DM}^{!}$ and $\textup{DM}^{*}$ can be computed via descent. Indeed, schemes locally of finite type admit open covers by affine schemes of finite type, and !!- and *-pullbacks agree up to shift and twist. The same reasoning shows that $\textup{DM}^{*}(X)\simeq\textup{DM}^{!}(X)$ for any algebraic space, since the diagonal of an algebraic space is representable by a scheme.
@@
Using étale descent, we may glue the six-functor formalism of DM to the restriction of $\textup{DM}^{*}$ to algebraic spaces, as in [Kha19], or equivalently via the DESCENT algorithm of Liu–Zheng [LZ17]. Similar constructions appear in [RS20], where ind-algebraic stacks and ind-schemes are treated.
@@
We may now apply the same arguments to glue the six-functor formalism to the restriction of $\textup{DM}^{*}$ to algebraic stacks locally of finite type over $S$. As before, the existence of a smooth cover implies that $\textup{DM}^{!}(X)\simeq\textup{DM}^{*}(X)$ for any such stack $X$.
To summarize, there is (up to equivalence) no distinction between working with $\textup{DM}^{!}$ and $\textup{DM}^{*}$. Furthermore, we obtain a complete six-functor formalism in the sense of Liu–Zheng for algebraic stacks locally of finite type over $S$, i.e., the functor $\textup{DM}^{*}$ extends to a symmetric monoidal functor
$\textup{DM}^{*}\colon\textup{Corr}({\mathrm{Stk}^{\textup{lft}}_{S}})_{\textup{lft}}\to\mathrm{CAlg}(\mathrm{Pr}^{L}_{\omega}).$
To simplify notation, we will henceforth write DM for the extension $\textup{DM}^{*}$. We retain the notation $\textup{DM}^{!}$ for the !!-variant.
@@
Recall the definition of a pro-$\mathcal{I}$}-algebraic stack of Definition 1. We simply say pro-algebraic if we do not want to specify $\mathcal{I}$}.
  1. $\bullet$
    If $(x,X)$ is a pro-$\mathcal{I}$}-algebraic stack, then we denote by $X_{i}\coloneqq x(i)$ and the transition maps $X_{j}\rightarrow X_{i}$ by $x_{ij}$, for $i\rightarrow j\in{\mathcal{I}}$.
  2. $\bullet$
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of pro-$\mathcal{I}$}-algebraic stacks. If $\bf P$} is a property of a morphism of algebraic stacks, we will say that $f$ has property $\bf P$} if each $f_{i}\colon X_{i}\rightarrow Y_{i}$ has property $\bf P$}.
  3. $\bullet$
    If $X$ is an algebraic stack, we will abuse notation to denote the constant diagram $i\mapsto X$ (here the transition maps are given by $\operatorname{id}_{X}$) with $X$ itself.
We will define some important properties of pro-algebraic stacks inspired from examples of the introduction.
@@ 3. $\bullet$
Let $(x,X)$ be a pro-$\mathcal{I}$}-algebraic stack. Then we say that $(x,X)$ is
@@
    tame if each $x_{ij}$ is smooth [^1] In practice the pro-algebraic stacks we consider are always tame. However, not every assertion in our construction will need this condition..
Remark 2.4.
The property strict will not be used heavily in this article. Nevertheless, we want to remark that the stack of displays by Lau admits a presentation as a strict pro-$\mathbb{N}$-algebraic stack. Perhaps a more straightforward example is the infinite dimensional affine space, via the projections $\mathbb{A}^{n+1}\rightarrow\mathbb{A}^{n}$.
As remarked in the introduction, we only get a $6$-functor formalism for certain classes of morphisms. Let us define these.
Definition 2.5.
If $f\colon(x,X)\rightarrow(y,Y)$ is a morphism of pro-$\mathcal{I}$}-algebraic stacks, then we call $f$ cartesian if each square
$X_{j}$$X_{i}$$Y_{j}$$Y_{i}$$\scriptstyle{x_{ij}}$$\scriptstyle{f_{j}}$$\scriptstyle{f_{i}}$$\scriptstyle{y_{ij}}$
is a pullback diagram. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of pro-$\mathcal{I}$}-algebraic stacks *-adjointable if the exchange morphism $f_{i}^{*}y_{ij*}\to x_{ij*}f_{j}^{*}$ induced by each square above is an equivalence and !!-adjointable if $f_{j!}x_{ij}^{!}\rightarrow y_{ij}^{!}f_{i!}$ is an equivalence.
The property cartesian is perhaps a more natural restriction for the existence of a $6$-functor formalism. Especially, when considering that homotopy invariance, T-stability and localization requires a good notion of the affine line, the projective space resp. taking complement in the category of pro-algebraic stacks. However, it turns out that the property adjointable is enough for just the existence of a $6$-functor formalism if $\mathcal{I}$$\simeq\mathbb{N}$. We will make this more precise late but let us give an example of adjointable morphisms of pro-algebraic stacks.
@@
  1. (1)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a cartesian morphism of pro-$\mathcal{I}$}-algebraic stacks, then $f$ is adjointable.
  2. (2)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of pro-$\mathcal{I}$}-algebraic stacks. If $y_{ij*}$ is fully faithful for all $i\to j\in{\mathcal{I}}$ (e.g. when $(y,Y)\cong S$) and $(x,X)$ is strict, then a simple computation shows that $f$ is *-adjointable.
    Indeed, we use that $\operatorname{id}\to x_{ij*}x^{*}_{ij}$ and $y_{ij}^{*}y_{ij*}\to\operatorname{id}$ are invertible to conclude that the morphism
    $f_{i}^{*}y_{ij*}\rightarrow x_{ij*}x^{*}_{ij}f_{i}^{*}y_{ij*}\simeq x_{ij*}f_{j}^{*}y_{ij}^{*}y_{ij*}\rightarrow x_{ij*}f_{j}^{*}$
    is also invertible.
    If $(x,X)$ and $(y,Y)$ in addition tame and $\operatorname{id}\to y_{ij}^{*}y_{ij\sharp}$ is invertible (again this holds if $(y,Y)\cong S$), then the same argument together with purity shows that $f$ is !!-adjointable.
@@
Let $L/K$ be an algebraic extension of fields. Let ${\mathcal{F}}^{L}_{K}$ be the filtered category of all finite field extensions $L/E/K$. Let ${\mathcal{F}}^{L}_{K}\rightarrow\textup{Fields}$ be the associated diagram in the category of fields. Then $\operatorname{Spec}(L)\cong\lim_{E\in{\mathcal{F}}^{L}_{K}}\operatorname{Spec}(E)$. Now let $\mathfrak{X}$} be an algebraic stack over $K$. For any $E\in{\mathcal{F}}^{L}_{K}$, we set ${\mathfrak{X}}_{E}\coloneqq{\mathfrak{X}}\times_{K}E$. Let $x_{E^{\prime}/E}\colon{\mathfrak{X}}_{E}\rightarrow{\mathfrak{X}}_{E^{\prime}}$ be the projection for a field extension $E^{\prime}/E$. Then $(x,{\mathfrak{X}}_{L})$ is a classical pro-${\mathcal{F}}^{L}_{K}$-algebraic stack. Naturally, we also get a classical pro-${\mathcal{F}}^{L}_{K}$-algebraic stack $(\iota,\operatorname{Spec}(L))$ defined by the inclusions of field extensions in ${\mathcal{F}}^{L}_{K}$. The associated morphism $(x,{\mathfrak{X}}_{L})\rightarrow(\iota,\operatorname{Spec}(L))$ is cartesian. Note that $(\iota,\operatorname{Spec}(L))$ and thus also $(x,{\mathfrak{X}}_{L})$ is not strict as for example any Galois extension $E\hookrightarrow E$ is a $\operatorname{Gal}(E^{\prime}/E)$-torsor.
The above can be generalized to the following case. We can replace $\operatorname{Spec}(K)$ by $S$ and assume that $S$ is connected. Let $\bar{s}$ be a geometric point of $S$. Then we can define $F_{S}$ as the category of pointed étale covers over $(S,\bar{s})$. The construction above can be repeated in this case for any algebraic stack $\mathfrak{X}$} over $S$.
@@
Let $(x,X)$ be a pro-$\mathcal{I}$}-algebraic stack. Then we define the $\infty$-category of rational motives on $(x,X)$ as